OpenStudy (anonymous):

Raul found that by increasing his speed by 5 km/h, he could drive 12 km in 0.2 hours. What was Raul's original speed?

OpenStudy (anonymous):

final speed=12/0.2=12/(1/5)=60 original speed = 60 -5 = 55km/h

OpenStudy (anonymous):

THANK YOU! ITS CORRECT

OpenStudy (anonymous):

yw